# IC-Custom: Diverse Image Customization via In-Context Learning

### v1

@@ Abstract
Image customization, a crucial technique for industrial media production, aims to generate content that is consistent with reference images. However, current approaches conventionally separate image customization into position-aware and position-free customization paradigms and lack a universal framework for diverse customization, limiting their applications across various scenarios. To overcome these limitations, we propose IC-Custom , a unified framework that seamlessly integrates position-aware and position-free image customization through in-context learning. IC-Custom concatenates reference images with target images to a polyptych, leveraging DiT’s multi-modal attention mechanism for fine-grained token-level interactions. We introduce the In-context Multi-Modal Attention (ICMA) mechanism with learnable task-oriented register tokens and boundary-aware positional embeddings to enable the model to correctly handle different task types and distinguish various inputs in polyptych configurations. To bridge the data gap, we carefully curated a high-quality dataset of 12k identity-consistent samples with 8k from real-world sources and 4k from high-quality synthetic data, avoiding the overly glossy and over-saturated synthetic appearance. IC-Custom supports various industrial applications, including try-on, accessory placement, furniture arrangement, and creative IP customization. Extensive evaluations on our proposed ProductBench and the publicly available DreamBench demonstrate that IC-Custom significantly outperforms community workflows, closed-source models like GPT-4o, and state-of-the-art open-source approaches. IC-Custom achieves approximately 73% higher human preference across identity consistency, harmonicity, and text alignment metrics, while training only 0.4% of the original model parameters.
@@ 1 Introduction
To enable effective training of our unified framework, we meticulously curated a high-quality dataset comprising 8,000 identity-consistent diptychs from real-world sources, addressing the critical limitation of existing methods that rely on synthetic data, which often produces artificial-looking results such as overly glossy textures or oversaturated colors. We further extend our approach to multi-reference scenarios by collecting 2,137 high-quality polyptych samples from real-world sources and synthesizing an additional 2,787 polyptych samples using our trained diptych model. This comprehensive multi-reference dataset enables our model to learn richer identity representations from distinct contexts and viewpoints.
@@ 3.2 Preliminaries and Background
[…]
where [,][,][ , ] denotes the concatenation operation along the sequence length dimension, $\boldsymbol{c_{T}}$ refers to the text tokens, $\boldsymbol{V}_{\theta}$ is the velocity field prediction network, the output of $\boldsymbol{V}_{\theta}$ is truncated to the sequence length of $\boldsymbol{Z}_{t}$. Despite remarkable advances, the conditional customization method do not support mask inputs, making it challenging to effectively adapt to position-aware customization applications.
@@ 3.3 In-Context Polyptych Customization
where [;][;][ ; ] denotes the concatenation operation along the channel dimension, $\tau$ denotes the tokenization operation, $\boldsymbol{X}_{0}^{mask}$ is the masked image diptych, $\boldsymbol{M}$ is the mask diptych, $\boldsymbol{V}_{\theta}$ represents velocity field predictor.
@@ 3.4 In-Context Multi-Modal Attention
where $RoPE(\cdot)$ denotes rotary position encoding [39], $Q_{vis}$, $K_{vis}$, $V_{vis}\in\mathbb{R}^{n\times d}$ are the query, key, and value matrices for visual tokens respectively, $Q_{txt}$, $K_{txt}$, $V_{txt}\mathbb{R}^{l\times d}$ are the query, key, and value matrices for text tokens respectively, $R_{task}^{i}\in\mathbb{R}^{4\times d}$ represents the $i$-th learnable task-oriented register tokens, and $\text{MHA}(\cdot)$ denotes the Multi-Head Attention operation. Our proposed ICMA module replaces all multi-modal attention in both the double blocks and single blocks of the original MM-DiT architecture.
@@ 3.5 In-Context Customization Data Curation
To address this challenge, we introduce ProductData , a curated high-quality customization dataset comprising nearly $8,000$ identity-consistent image pairs collected from e-commerce platforms, representing authentic real-world scenarios. This dataset encompasses various data types including clothing try-on, cosmetics, furniture, electronics, accessories, home decor, kitchenware, and personal care products, with image resolutions ranging from $800\times 800$ to $3000\times 3664$ pixels. Additionally, to complement our product-focused dataset and enhance diversity, we carefully filtered the SynCD $1024\times 1024$ subset [18], resulting in $2787$ high-quality identity-consistent paired synthetic samples that cover a broader range of subjects and scenarios beyond commercial products. This constitutes our final diptych dataset of 11,000 samples. For polyptych extension, we curated 2,137 real-world samples and synthesized 2,787 additional samples using our diptych model, providing sufficient diversity for multi-reference customization training.
@@ 4.1 Experiments Setup
IC-Custom is built upon a pre-trained text-to-image model FLUX.1-Fill-dev [20]. We train LoRA [12] with rank 64 on the first 10 layers of both single blocks and double blocks, and directly fine-tune the image input layer and text input layer. The total trainable parameters amount to $49.26$M, which is only $0.4\%$ of the original FLUX model’s $12$B parameters (consisting of 19 double blocks and 38 single blocks). Compared to previous methods [38, 26] that train LoRA on all layers (e.g., DreamO [26] trained 707M parameters), our approach significantly reduces training costs. The model is optimized on our proposed 12k dataset using the AdamW [24] optimizer with a learning rate of $55e-$ 5 and a batch size of 4. We train single-scene reference customization for 20k iterations in the first stage, followed by 10k iterations of multi-reference customization in the second stage.

### v2

@@ Abstract
Image customization, a crucial technique for industrial media production, aims to generate content that is consistent with reference images. However, current approaches conventionally separate image customization into position-aware and position-free customization paradigms and lack a universal framework for diverse customization, limiting their applications across various scenarios. To overcome these limitations, we propose IC-Custom , a unified framework that seamlessly integrates position-aware and position-free image customization through in-context learning. IC-Custom concatenates reference images with target images to a polyptych, leveraging DiT’s multi-modal attention mechanism for fine-grained token-level interactions. We introduce the In-context Multi-Modal Attention (ICMA) mechanism with learnable task-oriented register tokens and boundary-aware positional embeddings to enable the model to correctly handle different task types and distinguish various inputs in polyptych configurations. To bridge the data gap, we carefully curated a high-quality dataset of 16K identity-consistent samples with 10K from real-world sources and 6K from high-quality synthetic data, avoiding the overly glossy and over-saturated synthetic appearance. IC-Custom supports various industrial applications, including try-on, accessory placement, furniture arrangement, and creative IP customization. Extensive evaluations on our proposed ProductBench and the publicly available DreamBench demonstrate that IC-Custom significantly outperforms community workflows, closed-source models like GPT-4o, and state-of-the-art open-source approaches. IC-Custom achieves approximately 73% higher human preference across identity consistency, harmonicity, and text alignment metrics, while training only 0.4% of the original model parameters.
@@ 1 Introduction
To enable effective training of our unified framework, we curated a high-quality dataset consisting of both real-world and synthetic samples. Specifically, we collected 8K identity-consistent diptychs from real-world sources and an additional 4K synthetic diptychs, resulting in a total of 12K diptychs. For multi-reference scenarios, we further collected 2K real-world polyptychs and synthesized 2K polyptychs, yielding a total of 4K polyptychs. This comprehensive dataset enables our model to learn robust identity representations across diverse contexts and viewpoints, while also addressing the limitations of previous methods that overly rely on synthetic data and often produce artificial-looking results.
@@ 3.2 Preliminaries and Background
$\min_{\theta}\int_{0}^{1}\mathbb{E}\left[\left\|\left(\boldsymbol{Z}_{1}-\boldsymbol{Z}_{0}\right)-\boldsymbol{V}_{\theta}\left([\boldsymbol{Z}_{t},\boldsymbol{Z}_{ref}],\boldsymbol{c_{T}},t\right)\right\|^{2}\right]dt$ (1)
where $[,]$ denotes the concatenation operation along the sequence length dimension, $\boldsymbol{c_{T}}$ refers to the text tokens, $\boldsymbol{V}_{\theta}$ is the velocity field prediction network, the output of $\boldsymbol{V}_{\theta}$ is truncated to the sequence length of $\boldsymbol{Z}_{t}$. Despite remarkable advances, the conditional customization method do not support mask inputs, making it challenging to effectively adapt to position-aware customization applications.
@@ 3.3 In-Context Polyptych Customization
where $[;]$ denotes the concatenation operation along the channel dimension, $\tau$ denotes the tokenization operation, $\boldsymbol{X}_{0}^{mask}$ is the masked image diptych, $\boldsymbol{M}$ is the mask diptych, $\boldsymbol{V}_{\theta}$ represents velocity field predictor.
@@ 3.4 In-Context Multi-Modal Attention
where $RoPE(\cdot)$ denotes rotary position encoding [39], $Q_{vis}$, $K_{vis}$, $V_{vis}\in\mathbb{R}^{n\times d}$ are the query, key, and value matrices for visual tokens respectively, $Q_{txt}$, $K_{txt}$, $V_{txt}\in\mathbb{R}^{l\times d}$ are the query, key, and value matrices for text tokens respectively, $R_{task}^{i}\in\mathbb{R}^{4\times d}$ represents the $i$-th learnable task-oriented register tokens, and $\text{MHA}(\cdot)$ denotes the Multi-Head Attention operation. Our proposed ICMA module replaces all multi-modal attention in both the double blocks and single blocks of the original MM-DiT architecture.
@@ 3.5 In-Context Customization Data Curation
To address this challenge, we introduce ProductData , a high-quality customization dataset designed to ensure both authenticity and diversity. Specifically, we first collect nearly 8K identity-consistent image pairs from e-commerce platforms, covering real-world scenarios such as clothing try-on, cosmetics, furniture, electronics, accessories, home decor, kitchenware, and personal care products, with image resolutions ranging from $800\times 800$ to $3000\times 3664$ pixels. To further enrich the dataset and broaden its coverage beyond commercial products, we supplement these real-world samples with 4K high-quality identity-consistent synthetic pairs, carefully filtered from the SynCD $1024\times 1024$ subset [18]. This results in a comprehensive diptych dataset of 12K samples. For polyptych extension, we additionally curate 2K real-world polyptych samples and synthesize 2K more using our diptych model, thereby providing sufficient diversity for multi-reference customization training.
@@ 4.1 Experiments Setup
IC-Custom is built upon a pre-trained text-to-image model FLUX.1-Fill-dev [20]. We train LoRA [12] with rank 64 on the first 10 layers of both single blocks and double blocks, and directly fine-tune the image input layer and text input layer. The total trainable parameters amount to 49.2649.26M, which is only $0.4\%$ of the original FLUX model’s 1212B parameters (consisting of 19 double blocks and 38 single blocks). Compared to previous methods [38, 26] that train LoRA on all layers (e.g., DreamO [26] trained 707M parameters), our approach significantly reduces training costs. The model is optimized on our proposed 12k dataset using the AdamW [24] optimizer with a learning rate of 5e55e-5 and a batch size of 4. We train single-scene reference customization for 20k iterations in the first stage, followed by 10k iterations of multi-reference customization in the second stage.